OpenStudy (anonymous):

will medal. A field house a section where the seating can be arranged so the first row has 9 seats, the second row has 11 seats, the third row has 13 seats and so on.. if there is sufficient space for 26 rows in the section, how many seats are there in that entire section?

OpenStudy (anonymous):

a collection of marbles is set up into 5 rows. the first row has 125 marbles, and each row after that one has 4 less marbles than the preceding row. how many total marbles are there in the 3rd row

OpenStudy (danjs):

each row increases # of seats by 2 chairs, one on each end, forming a triangle

OpenStudy (danjs):

for example first 3 rows 9 + 11+13+15+17 .....(26 numbers to add up)

OpenStudy (danjs):

sum of the ODD numbers starting at 9, (26 numbers)

OpenStudy (danjs):

Odd number for row n, a(n) = 2*(n+3)+1 For n is the row number starting at n = 1, to n=26, need to sum them all

OpenStudy (danjs):

a(1) = 2*(4) + 1 = 9 a(2) = 2*(5) + 1 = 11 a(3) = 2*(6) + 1 = 13 ... ... ... a(26) = 2*(29) + 1 = 59

OpenStudy (danjs):

are you doing series in class?

OpenStudy (anonymous):

no, this a review from way back and i totally forget

OpenStudy (danjs):

use \(\large\color{blue}{a(n) = 2*(n+3)+1 }\) for n=1 up to n=26 and add all the numbers It is just the sum of 26 odd numbers starting with 9 9+11+13+15+17+ .......59

OpenStudy (anonymous):

so the answer would be 59?

OpenStudy (danjs):

no, the 26th row has 59 chairs. you have to add together 26 odd numbers, starting with 9 and up to 59

OpenStudy (danjs):

9 + 11 + 13 + 15 + 17 + 19 + 21 + 23 +...........up to + 59

OpenStudy (anonymous):

884

OpenStudy (danjs):

\[\sum_{n=1}^{26}[2*(n+3) + 1]\] = 884

OpenStudy (danjs):

yep that is what i got

OpenStudy (danjs):

\[\large\color{blue}{\sum_{n=1}^{26}2*(n+3)+1 = 884 }\]

OpenStudy (anonymous):

a collection of marbles is set up into 5 rows. the first row has 125 marbles, and each row after that one has 4 less marbles than the preceding row. how many total marbles are there in the 3rd row Help with this one?

OpenStudy (danjs):

First row 125 Second - 125 - 4 = 121 third - 121 - 4 = 117

OpenStudy (danjs):

125 - 4(n-1) n = row number

OpenStudy (danjs):

n=1 125 n=2 121 n=3 117

OpenStudy (anonymous):

i got 113

OpenStudy (danjs):

first row has 125 , each row is 4 less than the previous, second row 125-4 third row 125-4-4

OpenStudy (danjs):

First row has 125 second row has 121 Third row has 117 Fourth Row has 113 Fifth Row has 109